a man is four time as old as his son,in four years time he will be three time as old as his son.what are their ages now?
Answer by Cromlix(4381)

Man is 4 x as old as his son
M = 4S
In 4 years time he will be 3 x as old as his son.
M + 4 = 3(S + 4)
Replace M on right hand side
with 4s
4S + 4 = 3(S + 4)
4S + 4 = 3S + 12
4S - 3S = 12 - 4
S = 8
Ages now:
Son 12 years old
Man 36 years old
Hope this helps.
:-)